Write a system of two equations in two unknowns for each problem. Solve each system by substitution.
Ticket sales. Tickets for a concert were sold to adults for $3 and to students for $2. If the total receipts were $824 and twice as many adult tickets were sold , then how many of each were sold?

Answer by checkley75(3666)

3*2S+2*S=824
6S+2S=824
8S=824
S=824/8
S=103 NUMBER OF STUDENT TICKETS SOLD.
PROOF
2*103=206 ADULT TICKETS SOLD.
PROOF
3*206+2*103=824
618+206=824
824=824